OpenStudy (anonymous):

2(h-8)-h=h-16 solve for h

OpenStudy (anonymous):

0

OpenStudy (karatechopper):

2(h-8)-h=h-16 2h-16-h=h-16 h-16=h-16 h-16+16=h-16+16 h=h all real numbers

OpenStudy (anonymous):

thank you very much